# Five parameters are all you need

### v1

@@ Abstract
The standard cosmological model, with its six independent parameters, has proven remarkably successful in describing the evolution of the Universe. One of these parameters, the optical depth to reionization $\tau_{\mathrm{reio}}$, represents the scatterings that Cosmic Microwave Background (CMB) photons will experience after decoupling from the primordial plasma as the intergalactic medium transitions from neutral to ionized. $\tau_{\mathrm{reio}}$ depends on the neutral hydrogen fraction $x_{\mathrm{HI}}$, which, in turn, should theoretically depend on cosmology. We present a novel method to establish the missing link between cosmology and reionization timeline. We discover the timeline has a universal shape well described by the Gompertz mortality law, applicable to any cosmology within our simulated data. This enables us to map cosmology to reionization using symbolic regression and to treat $\tau_{\mathrm{reio}}$ as a derived parameter. Reanalyzing CMB data with our universal $x_{\mathrm{HI}}$ tightens the constraint on $\tau_{\mathrm{reio}}$ by more than one order of magnitude to $\approx 1\%$ and reduced the error on the amplitude of the primordial fluctuations by a factor of 2.5 compared to Planck’s PR3 constraint. While our results rely on the astrophysical assumptions in our simulations, the methodology itself is independent of these assumptions; after all, $\tau_{\mathrm{reio}}$ is fundamentally a function of cosmology.
@@
The $\Lambda$CDM cosmological model has proven extremely effective in predicting the evolution of our Universe, relying on only six parameters 1. In particular, it explains the transition from a predominantly neutral state in the early stages to the familiar ionized intergalactic medium (IGM) observed in our relatively nearby surroundings. This transition is known as cosmic reionization. Despite a comprehensive understanding of the astrophysical principles governing this transition, uncertainties persist regarding its precise timeline 2. The advent of the James Webb Space Telescope (JWST) 3 represents a pivotal moment, substantially bolstering our ability to directly constrain the evolution of the neutral hydrogen fraction $x_{\mathrm{HI}}$. This progress is being driven by the JWST’s enhanced detection capabilities, enabling the observation of high-redshift quasars 4 and high-redshift galaxies 5, 6, 7, 8.
@@
Reionization leads to scattering of Cosmic Microwave Background (CMB) photons by free electrons, disrupting the CMB angular power spectra ($C_{\ell}$). This scattering suppresses the signal at scales smaller than the Hubble scale at reionization (approximately $\ell>10$) 9 due to the optical depth $\tau_{\mathrm{reio}}$. Additionally, it introduces a new signal in the polarization of CMB photons at large angular scales 1, that is $\propto\tau_{\mathrm{reio}}$ in $C^{TE}_{\ell}$, the cross-correlation of the $E$-mode polarization with the temperature (intensity), and is $\propto\tau_{\mathrm{reio}}^{2}$ in $C^{EE}_{\ell}$, the $E$-mode polarization angular auto power spectrum. Consequently, heightened sensitivity to CMB polarization becomes crucial for mitigating the degeneracy between $\tau_{\mathrm{reio}}$ and other cosmological parameters, particularly $A_{\mathrm{s}}$, the amplitude of the primordial scalar power spectrum, and $r$, the ratio of tensor-to-scalar modes 10.
@@
Given the challenges posed by $\tau_{\mathrm{reio}}$ in CMB analyses and the anticipated advancements in constraining the reionization timeline 15, 16, now is an opportune moment to reassess its role. In theory, cosmic reionization is uniquely determined given a specific cosmology, i.e. by the other five cosmological parameters. Specifically, the evolution of the global fraction of neutral hydrogen can be written as
$x_{\mathrm{HI}}(a)=f(a;\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}})\quad\Rightarrow\quad\tau_{\mathrm{reio}}=g(\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}}),$ (1)
where $n_{\mathrm{s}}$, $h$, $\Omega_{\mathrm{b}}$, and $\Omega_{\mathrm{m}}$ are the tilt of the primordial power spectrum, dimensionless Hubble constant, and present baryon and matter density fractions, respectively. $\sigma_{8}$ is the present linear rms relative density fluctuation in a sphere of radius $8\,h^{-1}$pc. This parametrization chosen for convenience fully determines $x_{\mathrm{HI}}$, but the incomplete understanding of cosmic reionization obscures this mapping, necessitating the introduction of $\tau_{\mathrm{reio}}$ in CMB analyses. However, our understanding of the astrophysical processes governing reionization has significantly improved 17, 18, 19, 20 since the inclusion of $\tau_{\mathrm{reio}}$ became a standard practice. Ongoing and forthcoming observations promise to further our understanding and reduce inherent modeling uncertainties. Motivated by these developments, we use symbolic regression (SR) 21 to construct a mapping between cosmology and reionization timeline, aiming to demote $\tau_{\mathrm{reio}}$ from an independent to a derived cosmological parameter, and simultaneously tightening constraints in cosmological parameters.
@@
Eq. 1 introduces a novel avenue to constrain cosmology by examining the dependence of $x_{\mathrm{HI}}(z)$ on cosmological parameters. This mapping can enhance parameter constraints and shed light on reionization astrophysics. It also aids ongoing efforts in parametrizing cosmic reionization models 22, 23, 12 by including the cosmological dependence of $x_{\mathrm{HI}}$.
@@
Here, we present a universality in the neutral hydrogen time evolution, and derive through SR its dependence on cosmology on simulated reionization histories from 21cmFAST 19. We integrate this shape into CLASS 24, a popular Boltzmann solver for CMB analyses. We then evaluate the modified CLASS alongside Cobaya 25, a speed-aware sampler26, 27111With adaptive covariance learning and fast-dragging as in 28, enabling larger steps in slow parameters via intermediate transitions of fast parameters., showcasing its ability to recover parameter constraints from CMB data, including ‘TTTEEE’ + lensing likelihoods 29, 30 (see Figure 9). Finally, we demonstrate cosmological gains by computing $\tau_{\mathrm{reio}}$ as a derived parameter using our approach – Eq. 1 – compared to sampling over $\tau_{\mathrm{reio}}$ using the conventional $\tanh$ model 31. We summarize our strategy in Figure 2.
@@
Figure 2: Strategy to demote $\tau_{\mathrm{reio}}$ to derived parameter. a) Sobol sampling of $\bm{\theta}$ comprising 5 cosmological and 1 astrophysical parameters (see Figure 5 in Extended Data). b) Simulated $x_{\mathrm{HI}}$ timelines as a function of $\bm{\theta}$ and scale factor $a$. c) With symbolic regression, we optimize the mapping from $\bm{\theta}$ to the rescaling parameters that bring the universality. d) We model the universal shape (upper panel) as a composition of the Gompertz function and a low-degree polynomial (lower panel). e) Planck CMB data we re-analyze. f) We infer the parameter constraints using Monte Carlo Markov Chain (MCMC).
@@
To achieve our goal outlined in Eq. 1, we construct a universal shape for $x_{\mathrm{HI}}$ using 128 Sobol samples of 21cmFAST simulations (see Simulations and Figure 5 in the Extended Data). All $x_{\mathrm{HI}}(a)$ profiles share this shape, with differences between scenarios being mere translations and rescalings. Reionization causes $x_{\mathrm{HI}}$ to reduce from near 1 to effectively 0 via a sigmoid transition. The standard $\tanh$ function is symmetric in nature and not flexible enough to provide the early start and rapid completion suggested by reionization simulations 22, 32. The Gompertz curve, an asymmetric sigmoid function often used to analyze age-dependent human mortality 33, proves a good model for the survival of neutral hydrogen too. Its expected accelerated increase in mortality with age resembles the expectation for the percolation of ionized HII bubbles during the end stages of reionization.
@@
One way to uncover the universality is to view each $x_{\mathrm{HI}}$ scenario as a cumulative probability distribution (CDF) in $-\ln a$. With this insight, we can translate and rescale each timeline using the mean and variance of its corresponding probability density function (PDF), i.e. $-\mathrm{d}x_{\mathrm{HI}}/\mathrm{d}\ln a$, and discover the existence of a universal shape followed by all scenarios. Therefore, cosmology only impacts the translation and rescaling parameters of each timeline, not its shape.
@@
$\bm{c}$ =$=\{0,1,0.1503,0.04850,0.005261,0.0002182\},$
$\tilde{a}(a;\bm{\theta})$ $=\Bigl{[}\frac{a}{\alpha(\bm{\theta})}\Bigr{]}^{\beta},$ (4)
where $\bm{\theta}$ denotes 6 astrophysical and cosmological parameters, $\alpha(\bm{\theta})$ is the power-law pivot (or logarithmic translation), and $\beta=8.290$ is the rescaling tilt, which, according to our 21cmFAST simulations, appears constant. This lack of cosmological dependence may stem from 21cmFAST’s treatment of reionization astrophysics. See Helium reionization for specifics on implementing HeI and HeII reionization.
Before fully leveraging our formalism to extract the cosmological dependence in the rescaling of Eq. 2 and relaxing the need for $\tau_{\mathrm{reio}}$ in CMB analyses, we first implement the Gompertz shape with independent $\tau_{\mathrm{reio}}$ in CLASS and confirm its agreement with the conventional $\tanh$ model (gomp 0 and $\tanh$ 0 in Table 1). Using Planck PR3 likelihoods ‘TTTEEE’ 29 and CMB lensing 30, we sample typical cosmological parameters with Cobaya 25, including $\tau_{\mathrm{reio}}$. Given a proposal for $\tau_{\mathrm{reio}}$, we determine the corresponding reionization timeline using bisection by varying $\ln\alpha$ for gomp 0, while for $\tanh$ 0, the reionization midpoint $z_{\mathrm{re}}$ is the tuning parameter. The sampler runs until the Gelman-Rubin statistic 34 satisfies $R-1<0.2$ for the between-chain variance of the confidence intervals. We repeat this for $\tanh$ 0 and verify the agreement between the two models.
@@
Gompertz-polynomial-shaped reionization can reproduce standard CMB analyses. Now, we move to establish the connection between the universal shape for $x_{\mathrm{HI}}$ and the rescaling of a given reionization scenario. We refer to this model as gomp 1. This rescaling is naturally a function of cosmology. For example, a larger density of matter $\Omega_{\mathrm{m}}$ results in deeper potential wells, accelerating structure formation and increasing the number of ultraviolet photons driving the reionization process. We employ PySR, an SR package, to establish the cosmological dependence of the rescaling in Eq. 4.
@@
While PySR initially guided us towards the Gompertz curve when directly regressing $x_{\mathrm{HI}}$, the final analysis only uses it to regress the pivot and tilt instead. We fit their values jointly with the polynomial coefficients as described above, and feed them as labels to the genetic algorithm to find the best analytic expression (see Symbolic regression in Extended Data for our definition of best). Using PySR we derived the following pivot as a function of cosmology
@@
$\ln\alpha(\bm{\theta})=(h^{\Omega_{\mathrm{m}}}+\sigma_{8})(\Omega_{\mathrm{b}}-n_{\mathrm{s}}-\Omega_{\mathrm{m}}),$ (5)
which, interestingly, is independent of the astrophysical parameter $\zeta_{\mathrm{UV}}$ (see Simulations in Extended data). This ionization efficiency modulates the timing of reionization in our simulations by regulating the abundance of photons that escape into the IGM. Its absence in Eq. 5 suggests that it either plays a minor role or its contribution can be mimicked by the cosmological parameters.
@@
Eqs. 4 and 5 imply that higher values of $n_{\mathrm{s}}$ hasten reionization by boosting power on small scales, fostering a greater abundance of ionizing sources and earlier completion 15. Note that our 21cmFAST simulations assume that faint galaxies are the primary drivers of reionization. Similarly, larger $\Omega_{\mathrm{m}}$ and $\sigma_{8}$ primarily expedite reionization by enhancing structure formation. Surprisingly, Eq. 5 suggests that higher $\Omega_{\mathrm{b}}$ delays reionization, likely due to the increased abundance of HI in the intergalactic medium requiring more ionizing photons.
@@
We note that within the prior range of our 21cmFAST simulations (see Simulations) and their corresponding astrophysics of reionization, the mapping derived from SR is not unique. Additional details and results using an alternative mapping – gomp 2 – are presented in Alternative mapping of the Extended Data. Nonetheless, our results are robust and independent of the choice of mapping.
@@
We implement Eq. 5 in our Gompertz CLASS, which given the cosmological parameters determines the pivot value and consequently the reionization history, $\tau_{\mathrm{reio}}$, and corresponding CMB angular power spectra. This gomp 1 model eliminates the need to sample over $\tau_{\mathrm{reio}}$ (or $z_{\mathrm{re}}$), requiring only five cosmological parameters, e.g. $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}}\}$. We use Cobaya to re-analyze the same CMB data, showcasing the full strength of our approach.
@@
Figure 3: Analysis of CMB data with reionization as a function of cosmology. The green contours represent our results using the Gompertz reionization model with Eq. 5, eliminating the need to sample over any reionization parameter. The blue contours correspond to the results obtained using the conventional $\tanh$ model, while the relevant Planck constraints 1 are depicted with gray lines for reference.
@@
Figure 3 underscores the impact of our universally-shaped Gompertz reionization model. This approach tightens the constraint on the optical depth to $\approx 1\%$, a remarkable improvement compared to $>10\%$ with the $\tanh$ prescription. Furthermore, the constraint on $A_{\mathrm{s}}$ improves dramatically since the TT data is no longer significantly hampered by the degeneracy between $A_{\mathrm{s}}$ and $\tau_{\mathrm{reio}}$. The error on $A_{\mathrm{s}}$ decreases by an impressive factor of 2.5 compared to Planck’s results 1. Overall, we recover tighter constraints across the board compared to Planck. See Figures 8 and 1 in the Extended Data for details.
@@
[…]
Our findings for the timeline of reionization align with late reionization scenarios, which are supported by high-$z$ Lyman-$\alpha$ observations 35, 49. However, recent discoveries by JWST indicate the presence of massive, bright galaxies at early redshifts $z\sim 10$ 5, 6, 7. The presence of these early galaxies suggests a potential preference for brighter galaxies to drive reionization, a role that in our 21cmFAST simulations was attributed to a population of faint galaxies.
Furthermore, our results are influenced by the semi-numerical prescription employed by 21cmFAST to ionize the IGM, which, while efficient and swift, could bias our findings. Moreover, our exploration within the astrophysical framework of 21cmFAST has been limited to varying the ionization efficiency (see Simulations in the Extended Data). Therefore, a more comprehensive exploration is warranted to ensure, for instance, that the rescaling tilt $\beta$ remains a constant. Additionally, a valuable exercise to refine the inherent relationship between cosmological parameters and reionization history would involve using more realistic, albeit slower, reionization models. One such option is to use the THESAN simulations 18, which are hydrodynamical simulations incorporating radiative transfer.
@@
To establish the universality of our proposed Gompertz model for the neutral hydrogen reionization and its relationship with the cosmological parameters, we conducted 128 21cmFASTv3 simulations to generate the corresponding $x_{\mathrm{HI}}(z;\bm{\theta})$ profiles. Our parameter space include $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}},\zeta_{\mathrm{UV}}\}$, comprising five cosmological and one astrophysical parameters. The selection of $\sigma_{8}$ instead of $A_{\mathrm{s}}$ is imposed by the input requirements of 21cmFAST [^2] Alternatively, one could use 21cmFirstCLASS50 to avoid $\sigma_{8}$.. We use a scrambled Sobol sequence 51, 52 to sample quasi-uniformly in the $\bm{\theta}$-space, within the following ranges:

### v2

@@ Abstract
The standard cosmological model, with its six independent parameters, successfully describes our observable Universe. One of these parameters, the optical depth to reionization $\tau_{\mathrm{reio}}$, represents the scatterings that Cosmic Microwave Background (CMB) photons will experience after decoupling from the primordial plasma as the intergalactic medium transitions from neutral to ionized. $\tau_{\mathrm{reio}}$ depends on the neutral hydrogen fraction $x_{\mathrm{HI}}$, which, in turn, should theoretically depend on cosmology. We present a novel method to establish the missing link between cosmology and reionization timeline using symbolic regression. We discover the timeline has a universal shape well described by the Gompertz mortality law, applicable to any cosmology within our simulated data. Unlike the conventional tanh prescription, our model is asymmetric in time and a good fit to astrophysical constraints on $x_{\mathrm{HI}}$. By combining CMB with astrophysical data and marginalizing over astrophysics, we treat $\tau_{\mathrm{reio}}$ as a derived parameter, tightening its constraint to $<3\%$. This approach reduces the error on the amplitude of the primordial fluctuations by a factor of 2.3 compared to Planck’s PR3 constraint and provides a commanding constraint on the ionization efficiency $\zeta_{\mathrm{UV}}=26.9^{+2.1}_{-2.5}$. We expect further improvements in the near term as reionization constraints increase and our understanding of reionization advances.
@@
The $\Lambda$CDM cosmological model has proven extremely effective in predicting the evolution of our Universe, relying on only six parameters 1. In particular, it explains the transition from a predominantly neutral state in the early stages to the familiar ionized intergalactic medium (IGM) observed in our relatively nearby surroundings. This transition is known as cosmic reionization. Despite a comprehensive understanding of the astrophysical principles governing this transition, uncertainties persist regarding its precise timeline 2. The advent of the James Webb Space Telescope (JWST) 3 represents a pivotal moment, substantially bolstering our ability to directly constrain the evolution of the neutral hydrogen fraction $x_{\mathrm{HI}}$. This progress is being driven by JWST’s enhanced detection capabilities, enabling the observation of high-redshift quasars 4 and high-redshift galaxies 5, 6, 7, 8.
@@
Reionization leads to scattering of Cosmic Microwave Background (CMB) photons by free electrons, disrupting the CMB angular power spectra $C_{\ell}$. This scattering suppresses the signal at scales smaller than the Hubble scale at reionization (approximately $\ell>10$) 9 due to the optical depth $\tau_{\mathrm{reio}}$. Additionally, it introduces a new signal in the polarization of CMB photons at large angular scales 1, that is $\propto\tau_{\mathrm{reio}}$ in $C^{TE}_{\ell}$, the cross-correlation of the $E$-mode polarization with the temperature (intensity), and is $\propto\tau_{\mathrm{reio}}^{2}$ in $C^{EE}_{\ell}$, the $E$-mode polarization angular auto power spectrum. Consequently, heightened sensitivity to CMB polarization becomes crucial for mitigating the degeneracy between $\tau_{\mathrm{reio}}$ and other cosmological parameters, particularly $A_{\mathrm{s}}$, the amplitude of the primordial scalar power spectrum, and $r$, the ratio of tensor-to-scalar modes 10.
@@
Given the challenges posed by $\tau_{\mathrm{reio}}$ in CMB analyses and the anticipated advancements in constraining the reionization timeline 19, 20, now is an opportune moment to reassess its role. Theoretically, cosmic reionization is uniquely determined by cosmology, i.e. $x_{\mathrm{HI}}(z)$ is fully determined by the other five cosmological parameters. However, incomplete understanding of reionization obscures this mapping, necessitating the additional empirical parameter $\tau_{\mathrm{reio}}$ in CMB analyses. Since the inclusion of $\tau_{\mathrm{reio}}$ became a standard practice, our understanding of the astrophysical processes governing reionization has significantly improved 21, 22, 23, 24 and ongoing and forthcoming observations promise to reduce inherent modeling uncertainties.
@@
Motivated by these developments, we use symbolic regression (SR) 25 to construct a mapping between cosmology, astrophysics, and reionization timeline, aiming to demote $\tau_{\mathrm{reio}}$ from an independent to a derived cosmological parameter and simultaneously tightening constraints in cosmological and astrophysical parameters. This mapping can also shed light on reionization astrophysics and aid ongoing efforts in parametrizing reionization models 26, 27, 13 by including the cosmological dependence of $x_{\mathrm{HI}}$.
@@
Here, we present a universality in the neutral hydrogen time evolution related via a power-law transform, and derive through SR the dependence of power-law parameters on cosmology and astrophysics from simulated 21cmFAST 28, 23 reionization histories. We integrate this universally shaped reionization timeline into CLASS 29, a popular Boltzmann solver for CMB analyses. We then evaluate the modified CLASS alongside Cobaya 30, a speed-aware sampler 31, 32, showcasing its ability to recover parameter constraints from CMB data, including ‘TTTEEE’ + lensing likelihoods 33, 34 (see Figure 9). Finally, combining CMB and astrophysical data, we marginalize over reionization astrophysics to compute $\tau_{\mathrm{reio}}$ as a derived parameter using SR, quantifying the gains compared to sampling over $\tau_{\mathrm{reio}}$ utilizing the conventional $\tanh$ model 35. We summarize our strategy in Figure 2.
@@
Figure 2: Strategy to demote $\tau_{\mathrm{reio}}$ to derived parameter. a) Sobol sampling of $\bm{\theta}$ comprising 5 cosmological and 1 astrophysical parameters (see Figure 5 in Extended Data). b) Simulated $x_{\mathrm{HI}}$ timelines as a function of $\bm{\theta}$ and scale factor $a$. c) With symbolic regression, we optimize the mapping from $\bm{\theta}$ to the rescaling parameters that bring the universality. d) We model the universal shape (upper panel) as a composition of the Gompertz function and a low-degree polynomial (lower panel). e) Planck CMB data and $x_{\mathrm{HI}}$ data we analyze. f) We infer the parameter constraints using Monte Carlo Markov Chain (MCMC).
@@
We construct the universal $x_{\mathrm{HI}}$ shape using 256 Sobol samples of 21cmFAST simulations, varying 5 cosmological parameters and the astrophysical ionizing efficiency $\zeta_{\mathrm{UV}}$. The latter modulates the timing of reionization by regulating the abundance of photons that escape into the IGM (see Simulations and Figure 5 in the Extended Data). All $x_{\mathrm{HI}}(a)$ profiles share the same shape, with differences between scenarios being mere translations and rescalings in logarithmic scale factor $\ln a$. Reionization causes $x_{\mathrm{HI}}$ to reduce from near 1 to effectively 0 via a sigmoid transition. The standard $\tanh$ function is symmetric in nature and not flexible enough to provide the early start and rapid completion suggested by reionization simulations 26, 36 and favored by astrophysical data (see Figure 4). In contrast, the SR-inspired Gompertz curve, an asymmetric sigmoid function often used to analyze age-dependent human mortality 37, proves a good model for the survival of neutral hydrogen too. Its expected accelerated increase in mortality with age resembles the expectation for the percolation of ionized hydrogen bubbles during the end stages of reionization.
@@
One way to uncover the universality is to view each $x_{\mathrm{HI}}$ scenario as a cumulative probability distribution (CDF) in $-\ln a$. With this insight, we can translate and rescale each timeline using the mean and variance of its corresponding probability density function (PDF), i.e. $-\mathrm{d}x_{\mathrm{HI}}/\mathrm{d}\ln a$, and discover the existence of a universal shape followed by all scenarios. Therefore, cosmology and $\zeta_{\mathrm{UV}}$ only impacts the translation and rescaling parameters of each timeline, not its shape.
@@
$\bm{c}$ =$=\{0,1,0.1130,0.02600,0.0005491,-0.00006518\},$
$\tilde{a}(a;\bm{\theta})$ $=\Bigl{[}\frac{a}{\alpha(\bm{\theta})}\Bigr{]}^{\beta(\bm{\theta})},$ (3)
where $\bm{\theta}$ denotes 6 astrophysical and cosmological parameters, $\alpha(\bm{\theta})$ is the power-law pivot (or logarithmic translation), and $\beta(\bm{\theta})$ is the rescaling tilt. Their parameter dependences stem from 21cmFAST’s modeling of reionization astrophysics. Also see Helium reionization for specifics on implementing HeI and HeII reionization.
Before fully leveraging our methodology to extract the parameter dependence in the rescaling of Eq. 1 and relaxing the need for $\tau_{\mathrm{reio}}$ in CMB analyses, we first implement the Gompertz shape with independent $\tau_{\mathrm{reio}}$ in CLASS and confirm its agreement with the conventional $\tanh$ model (gomp and $\tanh$ in Table 1). Using Planck PR3 likelihoods ‘TTTEEE’ 33 and CMB lensing 34, we sample typical cosmological parameters with Cobaya 30, including $\tau_{\mathrm{reio}}$. Given a proposal for $\tau_{\mathrm{reio}}$, we determine the corresponding reionization timeline using bisection by varying $\ln\alpha$ for gomp while fixing $\beta$ to its mean value, meanwhile for $\tanh$, the reionization midpoint $z_{\mathrm{re}}$ is the tuning parameter. The sampler runs until the Gelman-Rubin statistic 38 satisfies $R-1<0.01$ for the variance in parameter means from different chains. We repeat this for $\tanh$ and verify the agreement between the two models.
@@
Having demonstrated that the universally shaped reionization can reproduce standard CMB analyses, we move to establish the connection between the universal shape for $x_{\mathrm{HI}}$ and the rescaling of a given reionization scenario. We refer to this model as gomp + SRFull. This rescaling naturally depends on cosmology and astrophysics. For example, a larger matter density $\Omega_{\mathrm{m}}$ results in deeper potential wells, accelerating structure formation and increasing the number of ultraviolet photons driving the reionization process. We employ PySR, an SR package, to establish the parameter dependences of the rescaling in Eq. 3.
@@
While PySR initially guided us towards the Gompertz curve when directly regressing $x_{\mathrm{HI}}$, the final analysis only uses it to regress the pivot and tilt instead. We fit their values jointly with the polynomial coefficients as described above, and feed them as labels to the genetic algorithm to find the best analytic expression (see Symbolic regression in Extended Data for our definition of best). Using PySR we derived the following mapping
@@
$\ln\alpha(\bm{\theta})$ $=\Bigl{(}\frac{\Omega_{\mathrm{b}}}{\Omega_{\mathrm{m}}}\Bigr{)}^{h}-(\sigma_{8}-0.04835)\bigl{(}n_{\mathrm{s}}+0.3558\ln(0.1123\zeta_{\mathrm{UV}})\bigr{)}-\Omega_{\mathrm{m}}-n_{\mathrm{s}},$ (4)
$\beta(\bm{\theta})$ $=\Bigl{(}\frac{0.005660^{\Omega_{\mathrm{m}}}}{0.6015}-\ln\bigl{(}\zeta_{\mathrm{UV}}-(\Omega_{\mathrm{m}}+n_{\mathrm{s}}h)^{15.05}\bigr{)}+h\Bigr{)}\ln{\Omega_{\mathrm{b}}}+\frac{h}{\sigma_{8}},$ (5)
where $n_{\mathrm{s}}$, $h$, $\Omega_{\mathrm{b}}$, and $\Omega_{\mathrm{m}}$ are the tilt of the primordial power spectrum, dimensionless Hubble constant, and present baryon and matter density fractions, respectively. $\sigma_{8}$ is the present linear rms relative density fluctuation in a sphere of radius $8\,h^{-1}$pc.
@@
Eqs. 3, 4 and 5 are analytic expressions and thus interpretable. For example, higher values of $\Omega_{\mathrm{m}}$, $\sigma_{8}$, and $\zeta_{\mathrm{UV}}$ hasten reionization by enhancing structure formation and increasing the abundance of ionizing photons. Similarly, larger $n_{\mathrm{s}}$ primarily expedites reionization by boosting power on small scales, leading to more ionizing sources and earlier completion 19. Keep in mind that our 21cmFAST simulations assume that faint galaxies are the primary drivers of reionization. Surprisingly, Eqs. 4 and 5 suggests that higher $\Omega_{\mathrm{b}}$ delays reionization, likely due to more HI in the intergalactic medium requiring additional ionizing photons. The ratio $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}$ in $\alpha$ supports this view. Moreover, increasing $h$ slightly delays reionization, consistent with an increase in physical baryon density and the expectation that galaxies and ionizing photons will be more spread out.
@@
We note that within the prior range of our 21cmFAST simulations (see Simulations) and their corresponding astrophysics of reionization, the mapping derived from SR is not unique. Additional details and results using an alternative mapping – SRHalf – are presented in Alternative mapping of the Extended Data. Nonetheless, our results are robust and independent of the choice of mapping.
@@
We implement Eqs. 4 and 5 in our Gompertz CLASS, which given the cosmological and astrophysical parameters, determines the pivot and tilt values, and consequently the reionization history, $\tau_{\mathrm{reio}}$, and CMB angular power spectra. This gomp + SRFull model eliminates the need to sample over $\tau_{\mathrm{reio}}$ (or $z_{\mathrm{re}}$), requiring only five cosmological parameters and $\zeta_{\mathrm{UV}}$. Moreover, thanks to SR, the model can constrain $\zeta_{\mathrm{UV}}$ using CMB and astrophysical data directly, a link that was not utilized in previous efforts 40. The advantage of sampling over $\zeta_{\mathrm{UV}}$ instead of $\tau_{\mathrm{reio}}$ is that powerful astrophysical constraints on $x_{\mathrm{HI}}$ can be used to marginalize over $\zeta_{\mathrm{UV}}$. We use Cobaya to re-analyze the same CMB data and include astrophysical data (see Astrophysical data of the Extended Data) to effectively constrain and marginalize over $\zeta_{\mathrm{UV}}$.
@@
Figure 3: Analysis of CMB and astrophysical data with reionization as a function of cosmology and astrophysics. The green contours represent our results using the Gompertz reionization model with Eqs. 4 and 5, which eliminates the need to sample over any conventional reionization parameter and uses quasar damping wing data on $x_{\mathrm{HI}}$ to constrain and marginalize over reionization astrophysics. The blue contours correspond to the results obtained using the conventional $\tanh$ model, while the relevant Planck constraints 1 are depicted with gray lines for reference. The $\tanh$ model performs poorly in fitting astrophysical data; thus, combining $\tanh$ with astrophysical data is ill-advised.
@@
Figure 3 underscores the impact of our universally-shaped Gompertz reionization model, tightening the optical depth constraint to $<3\%$ compared to $>10\%$ with the $\tanh$ prescription. Furthermore, the constraint on $A_{\mathrm{s}}$ improves dramatically since the TT data is no longer significantly hampered by the degeneracy between $A_{\mathrm{s}}$ and $\tau_{\mathrm{reio}}$. The error on $A_{\mathrm{s}}$ decreases by an impressive factor of 2.3 compared to Planck’s results 1. Overall, we recover tighter constraints across the board compared to Planck. Notably, combining SR with CMB and astrophysical data, specifically quasar damping wing (DW) data, yields a highly competitive constraint on $\zeta_{\mathrm{UV}}=26.9^{+2.1}_{-2.5}$, a commanding improvement over previous constraints (e.g. $\zeta_{\mathrm{UV}}=28^{+52}_{-18}$) 40. See MCMC inference, Figures 8 and 1 in the Extended Data for details.
Figure 4: Optical depth evolution $\tau(z)$ and reionization history $x_{\mathrm{HI}}(z)$. Our best-fit gomp + SRFull model (green dotted line) of Planck CMB data + quasar damping wing (DW) data is asymmetric and differs significantly from that of the symmetric $\tanh$ model (blue solid line). Note $z$ is shown in logarithmic scale of $a^{-1}$. We also include an alternative mapping from physical parameters to Gompertz timeline – gomp + SRHalf – in the purple dashed line (see Alternative mapping). The shaded regions in the upper panel correspond to the inferred range in $\tau_{\mathrm{reio}}$ from analyzing Planck PR3 data. Additionally, the lower panel includes observational constraints from high-redshift quasars and galaxies (see Astrophysical data).
@@
Our findings for the timeline of reionization align with late reionization scenarios, which are supported by high-$z$ Lyman-$\alpha$ observations 39, 41. However, recent discoveries by JWST indicate the presence of massive, bright galaxies at early redshifts $z\sim 10$ 5, 6, 7. The presence of these early galaxies suggests a potential preference for brighter galaxies to drive reionization, a role that in our 21cmFAST simulations was attributed to a population of faint galaxies.
We note that our results assume a flat Universe. Relaxing this assumption requires adding an extra parameter to the functional form of $\tau$ 42. Furthermore, our results are influenced by the semi-numerical prescription employed by 21cmFAST to ionize the IGM, which, while efficient and swift, could bias our findings. Moreover, our exploration within the astrophysical framework of 21cmFAST has been limited to varying the ionization efficiency (see Simulations in the Extended Data). Therefore, a more comprehensive exploration is warranted to ensure, for instance, that our findings for $\alpha$ and $\beta$ will not bias cosmological analyses when incorporating additional astrophysical parameters, such as the minimum halo mass that host ionizing sources. Additionally, a valuable exercise to refine the inherent relationship between cosmological parameters and reionization history would involve using more realistic, albeit slower, reionization models. One such option is to use the THESAN simulations 22, which are hydrodynamical simulations incorporating radiative transfer.
Throughout this work we considered only DW for astrophysical data, but including current luminosity function constraints can already halve the error on $\tau_{\mathrm{reio}}$. Hence, as reionization constraints improve, we anticipate substantial cosmological gains with our framework.
@@
To establish the universality of our proposed Gompertz model for the neutral hydrogen reionization and its relationship with the cosmological parameters, we conducted 256 21cmFASTv3 simulations to generate the corresponding $x_{\mathrm{HI}}(z;\bm{\theta})$ profiles. Our parameter space include $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}},\zeta_{\mathrm{UV}}\}$, comprising five cosmological and one astrophysical parameters. The selection of $\sigma_{8}$ instead of $A_{\mathrm{s}}$ is imposed by the input requirements of 21cmFAST [^1] Alternatively, one could use 21cmFirstCLASS43 to avoid $\sigma_{8}$.. We first use a scrambled Sobol sequence 44, 45 of length 128 to sample quasi-uniformly within the following $\bm{\theta}$-ranges: